THE MULTIMODE APPROACH TO CHALLENGING PROBLEMS IN VIBRATION SPECTROSCOPY

\underline{JOEL M. BOWMAN}, STUART CARTER, and XINCHUAN HUANG, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322; NICHOLAS HANDY, Dept. of Chemistry, University of Cambridge, Cambridge, UK.

I will briefly describe the code ”MULTIMODE”, developed in collaboration with Stuart Carter and Nicholas Handy, and then present applications to the series of molecules H_{3}O^{+}, H_{3}O_{2}^{−} and H_{5}O_{2}^{+}. I will also describe recent collateral work on developing full dimensional ab initio-based potential energy surfaces for these molecules. Some comparisons with Diffusion Monte Carlo calculations done by Anne McCoy on these molecules will also be presented.